### v2

# VGC-Bench: A Benchmark for Generalizing Across Diverse Team Strategies in Competitive Pokémon

Abstract
Developing AI agents that can robustly adapt to dramatically different strategic landscapes without retraining is a central challenge for multi-agent learning. Pokémon Video Game Championships (VGC) is a domain with an extraordinarily large space of possible team configurations of approximately – far larger than those of Dota or StarCraft. The highly discrete, combinatorial nature of team building in Pokémon VGC causes optimal strategies to shift dramatically depending on both the team being piloted and the opponent’s team, making generalization uniquely challenging. To advance research on this problem, we introduce VGC-Bench: a benchmark that provides critical infrastructure, standardizes evaluation protocols, and supplies human-play datasets and a range of baselines – from large-language-model agents and behavior cloning to reinforcement learning and empirical game-theoretic methods such as self-play, fictitious play, and double oracle. In the restricted setting where an agent is trained and evaluated on a single-team configuration, our methods are able to win against a professional VGC competitor. We extensively evaluate all baseline methods over progressively larger team sets and find that even the best-performing algorithm in the single-team setting struggles at scaling up as team size grows. Thus, policy generalization across diverse team strategies remains an open challenge for the community. Our code is open sourced at https://github.com/cameronangliss/VGC-Bench.
1 Introduction
Pokémon is the highest-grossing media franchise in the world, with an estimated total revenue exceeding $100 billion and a global player base numbering in the millions. In 2008, the Pokémon Company launched the Video Game Championships (VGC), a competitive series of tournaments featuring significant cash prizes and international prestige. The largest tournament to date, EUIC 2025, featured 1,257 competitors. Despite this popularity and competitive depth, to the best of our knowledge, no AI system has yet achieved human-level – let alone superhuman – performance in competitive Pokémon VGC battles.
Competitive Pokémon presents two tightly-coupled challenges: team building and team usage. We estimate the size of the possible team configuration space in VGC to be on the order of , vastly exceeding the contextual complexity of other benchmark games such as Dota or StarCraft. Because of the discrete and combinatorial nature of team-building, optimal strategies can vary dramatically depending on the team compositions of both the player and opponent. This results in extreme context sensitivity, where even expert human players struggle to generalize strategies across matchups – even when their own team remains fixed, as is standard in VGC tournaments, but especially if trying to pilot many different team compositions. Some teams focus on controlling the weather; some focus on controlling the speed of the opponent; some focus on trapping the opponent’s Pokémon on the field. There is no well-defined maximum number of strategic paradigms with which a team can be built around, and some teams can be hybrid, playing to more than one of these powerful strategies at the same time.
This combination of a vast space of team configurations and highly diverse team strategies makes Pokémon an especially valuable testbed for generalization in AI systems. In the past, research has been conducted on Pokémon with heuristic and search-based agents [14, 12, 10], reinforcement learning [9, 20, 25, 15], team building [16], and predicting the winner at any state [4].
In this work, we present VGC-Bench, a benchmark designed to evaluate AI generalization in Pokémon VGC. Our contributions include infrastructure for multi-agent learning and human-play data collection, a suite of competitive baselines, and robust evaluation tools for performance, generalization, exploitability, and human interaction. In the restricted setting where an agent is trained and evaluated on a single-team configuration, our methods are able to win against a professional VGC competitor. We then extend evaluation to broader multi-task settings involving diverse team configurations. While existing methods can learn competent policies in narrow conditions, they exhibit notable performance degradation as team diversity increases. These findings underscore the need for more generalizable learning algorithms in multi-agent, multi-task environments like VGC.
Furthermore, as part of our work, we contributed significant open-source improvements to Poke-env [17], a widely used library that has supported prior research in Pokémon AI. Our contributions include full integration with the PettingZoo [21] multi-agent framework and extended support for VGC and doubles formats. We expect that these enhancements will enable and encourage more accessible and rigorous experimentation, further advancing AI research in this domain.
2 Problem Formulation
VGC Mechanism: Pokémon VGC is a team-based competitive game with stochastic dynamics, simultaneous move selection, and a vast configuration space of possible teams. In the VGC format, each player assembles a team of 6 Pokémon, with each Pokémon configured by customized individual stats, up to 4 moves, a passive ability, a Tera type, and potentially a held item. VGC battles follow the doubles format, where players deploy up to 2 Pokémon on the field at a time. Recent tournaments adopt Open Team Sheets (OTS), a modification on the VGC format that reveals nearly all aspects about the opponent’s team – such as moves, items, abilities, and Tera types – while leaving the precise underlying stats concealed. Each match begins with a Team Preview phase, during which players simultaneously select 4 out of their 6 Pokémon to bring into the battle. The first two chosen are sent out at the start of the match, while the other two remain in reserve and can be switched in as the battle progresses. Once the battle begins, both players issue commands independently and simultaneously each turn, with Pokémon switching decision generally occurring before move choices. When all active Pokémon select moves, the execution order of actions is usually determined by their speed stats and speed ties are broken randomly. The objective is to knock out all of the opponent’s Pokémon by reducing their HP to zero before your own team is eliminated.
Several factors make VGC particularly challenging for AI agents:
1.
Combinatorial Team Configurations: The team configuration space in Pokémon VGC is extraordinarily large. With hundreds of Pokémon species, moves, items, and abilities, 19 possible Tera types per Pokémon, and a high-dimensional space of possible stat allocations, we estimate the total number of valid team configurations to be approximately . This combinatorial explosion far exceeds that of many other strategic games (see Table 1 for comparison).
2.
Stochastic Battle Mechanics: Move outcomes in VGC involve randomness in damage and secondary effects (e.g., 10% chance to paralyze). This results in a large branching factor per turn, which we approximate to be at times as large as .
3.
Partial Observability: Even with OTS, Pokémon VGC remains a partially observable game as the critical stats are not revealed. We estimate that the size of the information set – the set of all possible states given a partial observation of the game – is lower-bounded at approximately .
4.
Simultaneous and Multi-Agent Actions: Pokémon VGC is a simultaneous game that requires up to four Pokémon to make a decision at the same time. This game feature introduces non-stationarity and pose challenges to accurate credit assignment during policy optimization. For example, suppose both of a player’s Pokémon target the same opponent Pokémon, and one of them uses a move the target is immune to, but the other knocks the Pokémon out in one hit. In this case, assuming the AI only observes the game state before and after the turn, the AI could never know which Pokémon caused the knock out.
5.
Team Preview Strategy and Generalization: Pokémon VGC features a unique challenge: only four of the six team members are used in each battle. This introduces a mini team-selection problem, where the player must pick the most effective subteam against the opponent’s lineup. This single decision can drastically alter the dynamics of the battle, and with a total of possible team preview decisions for each player, effective exploration and generalization are especially challenging.
Formalization.
We model each Pokémon match as a two-player zero-sum partially observable stochastic game (POSG) with randomized team configurations. Let denote the finite set of legal team configurations, and let be the uniform distribution over configurations. At the beginning of each episode, players independently sample their teams , which jointly determine the instantiation of a POSG . Here, denotes the full battle state (including HP, status, weather, stat boosts, etc.), is the partial observation available to player , and is the set of available actions (select moves or switch Pokémon). is the stochastic transition function governed by Pokémon mechanics, and is the terminal reward function defined as if player 1 wins in terminal state and if player 2 wins. denotes the team configuration for both players.
Given policies and for the two players that map observation histories to action distributions, the expected return for player 1 across all team match-ups is defined as
where is the terminal state of trajectory . Since the game is zero-sum, the benchmark seeks a Nash equilibrium in expectation over configurations, which corresponds to the minimax-optimal pair of policies :
This formalism captures our objective: to find robust policies that perform well on average against all possible opposing team draws in the configuration space.
3 Related Work
Progress in Game AI.
Early AI milestones focused on perfect-information games like Chess and Go, where IBM’s Deep Blue and DeepMind’s AlphaGo achieved superhuman performance through massive search and pattern recognition [3, 19]. IBM’s Watson extended AI capabilities to natural-language trivia by winning Jeopardy! [5]. Poker introduced reasoning under partial observability and uncertainty, with Libratus defeating top humans in no-limit Texas Hold’em [2]. Recent work shifted to more complex multi-agent games: AlphaStar reached grandmaster in StarCraft II by population based reinforcement learning [24] and OpenAI Five mastered Dota 2 with limited hero drafting [1]. In contrast, Pokémon VGC offers a unique challenge that combines partial observability, stochasticity, simultaneous actions, and a combinatorially vast team configuration space larger than any prior benchmarked game (see Table 1).
| Game | Init. Config. Space | Partial Obs. | Stochastic | Simul. Actions |
| Chess | 2 | |||
| Go | 2 | |||
| Jeopardy! | 1 | |||
| Poker | ✓ | ✓ | ||
| StarCraft II | ✓ | ✓ | ✓ | |
| Dota 2 | ✓ | ✓ | ✓ | |
| Pokémon VGC | ✓ | ✓ | ✓ |
Existing Pokémon AI.
Early Pokémon AI research focused on handcrafted rules and search-based methods, such as those featured in the Showdown AI Competition[12], with agents using expectimax or minimax search[9], and heuristic bots designed to simulate human play [14]. Karten et al. [10] advanced this line with PokeChamp, combining minimax with LLMs for action sampling and opponent modeling. Deep reinforcement learning approaches have since demonstrated strong performance [8]: PPO agents trained via self-play[9], using Poke-env [20], or augmented with MCTS [25] have achieved expert-level play. However, most of the prior work focuses exclusively on the Single Battle format, where each player sends out one Pokémon at a time. This setting significantly simplifies the underlying decision space: in contrast, the official Pokémon VGC tournaments use Double Battle format, which involves two active Pokémon per side, leading to a combinatorial explosion in possible actions and interactions and more challenging learning. Moreover, prior evaluations are largely restricted to the Pokémon Showdown ladder, an online platform where opponents are often casual players of widely varying skill levels; by comparison, VGC-bench agents are directly evaluated against human experts and professional VGC players.
4 VGC-Bench
This section introduces VGC-Bench, a comprehensive benchmark built on top of poke-env [17]. VGC-Bench provides users with (1) infrastructure for multi-agent learning and human-play data collection, (2) a diverse suite of baselines, and (3) robust evaluation tools for performance, generalization, exploitability, and human play.
4.1 Infrastructure
Multi-Agent Environment.
We integrate poke-env with PettingZoo [21] to support parallelized population-based and PSRO[11]-style multi-agent learning for both players and support the VGC format. To support more controlled training settings, we introduce two toggles: one to exclude the team preview phase and another to disable mirror matches. (1) Disabling team preview skips that stage from the agent’s perspective, initializing each game with a random team preview decision – potentially encouraging policy exploration in any team preview situation. (2) Disabling mirror matches prevents an agent from facing the same team configuration during training, which is useful when evaluating a specific matchup between two distinct team configurations. However, we found that excluding the mirror match can result in reward hacking, where the agent learns to intentionally lose when playing one team and exploit the matchup when playing the other. Finally, we provide users with a pool of high-quality teams often used in real VGC competitions from VGCPastes. Teams are supplied as part of the environment configuration as plain text when a battle starts.
Observation Space.
The default observation encodes information for each of the 12 Pokémon (6 per player) using a mixture of discrete (e.g., Pokémon type) and continuous (e.g., HP) features. Let be the global features (e.g., weather), the side-specific features (e.g., light screen), and the per-Pokémon features. Each Pokémon is represented as a concatenation of , its side’s , and the global vector, giving a vector of size . The overall observation is shaped as . If frame stacking is used with frames, the observation becomes .
Action Space.
We represent actions in the VGC format as a joint action space, since each turn in doubles matches requires an action for each of the two active Pokémon. The action space is enumerated in Table 9. Each Pokémon has 107 available actions, which captures the full space of switching in benched Pokémon and using moves, where moves involve which move is being picked, the intended target, and whether or not the Pokémon is terastallizing (or using another available gimmick). We also unify the team preview action of the battle with the action space for the rest of the battle by simply modeling team preview as two joint "switch-in" actions in a row, providing two Pokémon per joint action for a total of four Pokémon to make the team preview decision.
Human-Play Data Collector.
To facilitate large-scale imitation learning from human players, we provide a parallel data collection pipeline for scraping and parsing Gen 9 VGC battle replays from Pokémon Showdown. With the current number of available replays from Pokémon Showdown, we are able to amass over 330,000 OTS-enabled VGC games. Due to the neutral perspective of battle logs and its purely transitional messages, reconstructing a trajectory of game states cannot be done perfectly. However, because we filter for only OTS-enabled battle logs, we can derive well-approximated states and actions from the logs by replaying the transitions through our environment.
4.2 Baseline Implementations
VGC-bench implements a total of 11 baseline agents: three heuristic agents, an LLM agent, a behavior cloning model (BC), a self-play RL agent (SP), a fictitious play RL agent (FP), and a double oracle RL agent (DO), as well as fine-tuning the agent initialized with the BC policy’s parameters using SP, FP, and DO (BCSP, BCFP, BCDO).
4.2.1 Trained Baselines
Policy and Value Network Architecture.
All the policy networks used across the learning baselines embed each agent’s moves, items, and abilities into latent representations, and then use an additional aggregation token and a 3-layer Transformer encoder [23] to aggregate the information of 12 Pokémon (6 from the ego teams and 6 from the opponent’s team) in the field. When frame stacking is enabled, we use a second Transformer encoder along the time axis, apply a positional encoding, and use causal masking to process the historical information and only use the logits from the last frame. The output of the Transformer encoder is then projected to the size of the action space by a linear layer into the logits before the softmax layer. The logits at the invalid actions are masked by . We also handle interdependent action constraints, such as ensuring that both Pokémon do not switch into the same replacement. For reinforcement learning methods, we adopt an actor-critic implementation without parameter sharing, though they share the same network architecture.
Behavior Cloning.
We train a reward-filtered behavior cloning (BC) policy [6] to match the distribution of human actions given a state from the dataset collected by the human play data collector. Specifically, we select trajectories from replays in which players have ladder ratings greater than or equal to , ensuring high-quality demonstrations. The parameterized policy is trained to imitate human decision-making by minimizing the negative log-likelihood of the demonstrated actions:
Population-based Reinforcement Learning.
We employ an actor-critic implementation of the Proximal Policy Optimization (PPO) [18] to train all reinforcement learning agents. For agents to approach the desired minimax solution of the game via RL, we implement three population-based training paradigms from the unified empirical game-theoretic framework [11] on top of RL: self-play, fictitious play, and double oracle [13]. In self-play (SP), the agent trains against itself. The fictitious play (FP) variant maintains a pool of the agent’s past checkpoints, and the agent learns against a uniform distribution of its past selves, resetting the opponent policies each episode. We take the final response policy to the policy pool as the method’s output. Double oracle (DO) differs from FP in that it derives the Nash equilibrium distribution from a maintained empirical payoff matrix between all agents in the pool, and uses that distribution to sample opponents. We solve the Nash distribution by solving a linear programming problem because of the game’s two-player, zero-sum, and symmetric-payoff nature. Future work could employ more complex meta-game solvers with the Policy Space Response Oracle [11].
Our preliminary experiments suggest that this population-based framework can potentially perform better when initialized with strong policies. Thus, we also implement baselines that fine-tune the policy from the BC-trained parameters (BCSP, BCFP, BCDO). Each method is trained with the same number of training teams and interaction timesteps. The hyperparameters for RL (details in Table 10) were tuned to serve as a reasonable starting point for future work.
4.2.2 Other Baselines
LLM Agent.
We also provide a basic LLM agent that uses Meta-Llama-3.1-8B-Instruct [7]. Each turn, the battle state is converted into a prompt that is then fed into the agent along with an available action space, instructions on how to construct a valid action, and the required format of its response. If the LLM agent does not respond in the correct format, the agent selects a random action. Note that the LLM agent was not thoroughly researched; we acknowledge that it is possible to make a stronger LLM agent with tool use and other techniques [10]. However, this LLM agent baseline provides a reasonable starting point on which research can be conducted.
Heuristic Agents.
From poke-env, we inherit 3 rule-based baselines: RandomPlayer, MaxBasePowerPlayer, and SimpleHeuristicsPlayer [17]. The RandomPlayer plays random moves every turn, the MaxBasePowerPlayer greedily chooses the highest-power move from the available options every turn, and the SimpleHeuristicsPlayer has a hard-coded weighted sum of multiple heuristics to select the best move based on the heuristic scores, based on information such as current HP of the active Pokémon, type matchup viability, accuracy, and power of its moves, etc. We extend SimpleHeuristicsPlayer so that it can also play VGC battles in a double format.
4.3 Evaluation Methods
VGC-Bench provides multiple methods of evaluation, including cross-evaluation of agents, training agents to exploit a fixed opponent, and a means to compute custom ELO ratings (the details of which are described in Appendix B). The cross-evaluation over games of a pair of policies from the policy pool , where , , and , can be defined point-wise as
where is a set of team configurations that the pair of policies both sample uniformly from each game. The combination of all these evaluation primitives enables insight into agent performance, generalization, and exploitability.
1.
Performance can be tested by cross-playing agents with teams that every agent had experience with during training. Ideally, as you train agents with more teams, performance would minimally decrease on an individual team basis compared with agents trained with fewer teams. We define as the set of teams configurations that policy was trained on; the evaluation set of teams is constrained by
(4) -
2.
Generalization can be tested by cross-playing agents with teams outside of their training distribution; if an agent trained with more teams outperforms agents trained on fewer teams, this suggests that the agent has generalization capabilities. We constrain the evaluation set of teams by
(5) -
3.
Exploitability can be tested by approximating a best response policy against the agent to be evaluated, , via RL training and measuring the highest win rate that the best response can achieve [22]. This provides a lower bound on the true exploitability of the agent. Whatever maximum win rate the counter-policy achieves over a sufficiently large number of training steps serves as our exploitability measurement for our agent. For an agent trained on teams, a best response policy must be trained for each individual team, and the exploitability of the agent is defined as
(6) In practice, we initialize our counter-policy with the BC agent since a random counter-policy initialization can struggle to get a reward signal against a strong agent. If the best response policy does not achieve a 100% win rate against the agent being tested, the user may run multiple trials to tighten the lower bound of the highest possible win rate that is being measured.
5 Experiments
To encourage research in Pokémon VGC, we gather preliminary results interrogating the following questions: (1) Is it possible to use standard RL methods to create an AI agent that plays at a human expert level for the one-team setting? (2) How do performance, generalization, and exploitability scale as the training team size increases? All experiments are conducted on a cluster with 8 A40 GPUs and 2 Intel(R) Xeon(R) Gold 6342 CPU @ 2.80GHz.
5.1 Cross-Play Performance
For all agents that fine-tuned with RL, we did training runs with team set sizes of 1, 3, 10, and 30. For each team set size, we compare our 6 RL baseline agents, the BC agent, 3 rule-based players, and our LLM agent (all described in Section 4.2.1) through cross-play in Figure 3. Each win rate entry of the cross-play matrix is calculated with episodes with a uniformly drawn team for each game played. For the actual win count values, see Tables 5, 6, 7, 8 in the Appendix C.
Notice that in the one-team setting (whose win counts are recorded in table 5), FP beats BCFP 83 to 17, BCFP beats DO 83 to 17, and DO beats FP 62 to 38 – evidence that Pokémon VGC is a cyclic game. However, the heatmaps become more transitive as the training team size increases.
We also calculate ELO ratings based on the payoff matrices for each team set size as shown in Table 2. Note that these ELO ratings are calculated within the VGC-Bench benchmark, not on the online pokemon-showdown ladder. The online ladder is not useful for evaluating performance on in-distribution teams since the opponent’s team configuration cannot be controlled, but we acknowledge that the ladder is a useful metric to measure out-of-distribution performance. It is clear from Figure 3 and Table 2 that while non-BC methods achieve decent performance, they are generally outperformed by methods initialized with BC. It is not clear from these results which of self-play, fictitious play, and double oracle are most effective, but fictitious play seems to be the most reliably high-performance.
| R | MBP | SH | LLM | SP | FP | DO | BC | BCSP | BCFP | BCDO | |
| 1 | 1127 | 1520 | 1621 | 1204 | 1681 | 1717 | 1573 | 1463 | 1721 | 1768 | 1627 |
| 3 | 1122 | 1443 | 1585 | 1174 | 1631 | 1609 | 1523 | 1370 | 1736 | 1689 | 1708 |
| 10 | 1124 | 1466 | 1602 | 1170 | 1634 | 1604 | 1562 | 1444 | 1692 | 1730 | 1743 |
| 30 | 1052 | 1398 | 1544 | 1069 | 1508 | 1538 | 1514 | 1361 | 1612 | 1647 | 1642 |
5.2 Performance and Generalization Evaluations
We then measure performance and generalization as described in Section 4.3. We used the highest overall ELO-rated agents in the 1, 3, 10, and 30-team settings for this evaluation; Table 4 shows that as the size of the set of teams used during training increases, the performance of the AI agent on any one team decreases considerably and consistently. These results indicate that our agent does not pass the performance test even for relatively small , highlighting the central challenge that Pokémon VGC poses to the AI community. To test generalization, we used 4 out-of-distribution teams and compared agents across team set sizes via cross-play. In Table 4, we can see that our agent exhibits generalization since win rate increases moderately as increases.
| #Teams | 1 | 3 | 10 | 30 |
| 1 | – | 79 | 83 | 92 |
| 3 | 21 | – | 68 | 82 |
| 10 | 17 | 32 | – | 60 |
| 30 | 8 | 18 | 40 | – |
| #Teams | 1 | 3 | 10 | 30 |
| 1 | – | 55 | 34 | 33 |
| 3 | 45 | – | 30 | 33 |
| 10 | 66 | 70 | – | 55 |
| 30 | 67 | 67 | 45 | – |
5.3 Exploitability Evaluation
We measured the exploitability of the strongest ELO-rated agent from each of our 4 team set sizes as described in Section 4.3. For each, we initialized the exploiter agent with the BC agent’s parameters. We found that for all 4 agents, after 3 million training steps, the exploiter agent was able to achieve approximately 100% win rate.
5.4 Human Evaluation
After measuring ELO ratings and evaluating performance across team set sizes, we tested our most performant agent in the most performant team set size – the behavior cloning agent fine-tuned with fictitious play in the 1-team setting – against an intermediate, advanced, and expert player. Against the intermediate player, our agent won all 5 matches, and against the advanced player, the agent won 2 out of 5 total matches played. Indeed, our agent was even able to win matches against the expert-level player who has qualified for the VGC World Championships. To the best of our knowledge, our agent is the first AI system to achieve such a feat. We received feedback that although the agent is very strong on initial play, it does have noticeable dips in performance in certain states. After enough successive games, strong human players can adapt and beat the agent.
6 Conclusion
In this work, we introduced VGC-Bench, a benchmark designed to evaluate the generalization capabilities of AI agents in the challenging and combinatorially complex environment of Pokémon VGC. Our benchmark includes standardized evaluation protocols, a modular training pipeline, curated human gameplay data, and implementations of a broad set of baseline methods ranging from behavior cloning and reinforcement learning to game-theoretic algorithms like fictitious play and double oracle. We also contributed substantial improvements to the poke-env library, including integration with PettingZoo and environment support for VGC formats, thereby enabling easy adoption by the broader research community.
Through extensive experiments, we demonstrated that while current algorithms can attain strong performance in the single-team setting, they degrade significantly when scaling to multi-team generalization – highlighting a key open challenge in multi-agent learning. By providing a reproducible benchmark and uncovering a difficult generalization frontier, our work establishes a foundation for future progress in robust multi-agent policy learning.
Looking forward, we identify two major research directions enabled by VGC-Bench:
1.
Generalization to Teams, : Our current agents achieve expert-level performance in the single-team setting but struggle as the number of evaluated teams increases. A natural extension is to develop agents that can generalize across team matchups, evaluate metagames through cross-play matrices, and ultimately perform at a superhuman level regardless of the team piloted. Researchers can use the experiments in Tables 4 and 4 to assess progress on this front.
2.
Team Building: A strong VGC agent could be used to evaluate candidate teams and provide a reward signal for searching the vast team configuration space. While such work would be downstream of generalization, our current agents may already offer a useful starting point for this line of research.
Appendix A Game Analysis
Pokémon VGC is an incredibly complex game; there are up to four Pokémon out at a time, each able to target up to three targets with a move and potentially terastallize or switch out into up to two benched Pokémon. Then we can calculate the approximate branching factor of nondeterminism in Pokémon moves. Calculated loosely, there are 16 damage rolls for each damaging move, almost always with a chance to hit normally, fail, land a critical hit, land an additional effect, or both land a critical hit and an additional effect. If we only consider these possibilities, there are outcomes for every move used. Also, each Pokémon has 4 total moves they can use, each of which have a maximum of 3 possible targets. If terastallization is available, there are
possible outcomes, and without terastallization there are
possible outcomes. Thus, in a turn, there is a worst-case branching factor of approximately
Additionally, we can quantify the worst-case size of a player’s information set, or the number of indistinguishable states due to Pokémon VGC’s partial observability. The only unobservables for a player are the opponent Pokémon’s stats and the Pokémon that haven’t yet been revealed to be brought by the opponent during team preview. For the latter, the opponent immediately reveals half of their team preview decision at the beginning of the battle, leaving only two unrevealed Pokémon left out of 4 possibilities. That leaves possibilities. Pokémon stats are determined by a Pokémon’s EV’s, IV’s, and natures. There are 25 natures (although 5 are identical in effect, so we will use 21), 32 possible values for each of its 6 IV values, and 512 points to distribute over 6 stats for EV values, with no more than 252 points for each stat. Assuming all EV points are used and all EV spreads would distribute points to stats that are divisible by 4, which is reasonable because EV points only affect stats in multiples of 4, we can formalize EV spreading as the problem of distributing units across 6 stats, with the maximum allocation per stat being . We must find the total number of solutions to the equation , for . This can be solved using the inclusion-exclusion principle for bounded integer compositions, and yields a total count of 246,774,528. Additionally, players almost always set all of their Pokémon’s IVs to 31, so we will leave the IV configuration space out of the calculation so as to not overestimate the realistic size of the configuration space. Thus, the number of possible realistic stat configurations is
This is the best-case size of the information set if the opponent has revealed all four Pokémon they chose during team preview; if they haven’t, then the worst-case size of the information set is
Let us now attempt a rough count of the total number of possible team configurations in VGC in Pokémon Scarlet and Violet. With 937 total moves in the game (of which we will approximate about 100 are available in a given Pokémon’s learnset), up to 3 available abilities for a species, 540 holdable items (of which only 223 are available in the latest game and are not pokeballs), 2 genders (usually), as well as the choice of a tera type and stat configurations for each Pokémon (calculated above to be ), the total number of configurations for a single Pokémon can be calculated as
Thus, with approximately 750 species available in the current games as of the writing of this paper, the total number of team configurations is
We will now compare this with the initial configuration space of other games. In chess and Go, players can only initialize as white or black, and in Jeopardy there is no initial configuration at all. In Texas hold ’em Poker, each player is dealt 2 cards out of a 52-card deck at the beginning of the game, yielding a total initialization space of for the first player, for the second player, etc., or a total configuration space of
Dota has two teams of five players choose from a total of 126 available heroes, yielding a total configuration space of . Finally, Starcraft allows each of the two players to choose from three available races, yielding total configurations.
Appendix B ELO Ratings
To calculate ELO ratings from a matrix of win rates for , we define the implied rating difference via the formula
so that
We then form the linear system
for all pairs with valid comparisons, and add the constraint
to fix the mean rating. We solve this overdetermined system in the least-squares sense to obtain preliminary ratings . Finally, we shift all ratings so that
and, if desired, rescale so that the standard deviation satisfies
Appendix C Numerical Cross-Play Results
| Algorithm | R | MBP | SH | LLM | SP | FP | DO | BC | BCSP | BCFP | BCDO |
| R | – | 6 | 1 | 38 | 1 | 2 | 4 | 5 | 0 | 1 | 4 |
| MBP | 94 | – | 40 | 88 | 29 | 48 | 33 | 53 | 22 | 6 | 27 |
| SH | 99 | 60 | – | 91 | 52 | 40 | 100 | 65 | 25 | 22 | 62 |
| LLM | 62 | 12 | 9 | – | 5 | 2 | 2 | 12 | 1 | 0 | 05 |
| SP | 99 | 71 | 48 | 95 | – | 30 | 95 | 73 | 20 | 61 | 62 |
| FP | 98 | 52 | 60 | 98 | 70 | – | 38 | 88 | 63 | 83 | 78 |
| DO | 96 | 67 | 0 | 98 | 05 | 62 | – | 83 | 40 | 17 | 08 |
| BC | 95 | 47 | 35 | 88 | 27 | 12 | 17 | – | 13 | 04 | 18 |
| BCSP | 100 | 78 | 75 | 99 | 80 | 37 | 60 | 87 | – | 38 | 55 |
| BCFP | 99 | 94 | 78 | 100 | 39 | 17 | 83 | 96 | 62 | – | 89 |
| BCDO | 96 | 73 | 38 | 95 | 38 | 22 | 92 | 82 | 45 | 11 | – |
| Algorithm | R | MBP | SH | LLM | SP | FP | DO | BC | BCSP | BCFP | BCDO |
| R | – | 9 | 1 | 48 | 3 | 2 | 4 | 16 | 0 | 2 | 0 |
| MBP | 91 | – | 38 | 85 | 25 | 21 | 35 | 66 | 8 | 12 | 10 |
| SH | 99 | 62 | – | 97 | 35 | 45 | 53 | 79 | 34 | 32 | 21 |
| LLM | 52 | 15 | 3 | – | 4 | 2 | 7 | 16 | 3 | 2 | 0 |
| SP | 97 | 75 | 65 | 96 | – | 53 | 69 | 90 | 38 | 46 | 35 |
| FP | 98 | 79 | 55 | 98 | 47 | – | 53 | 87 | 17 | 38 | 33 |
| DO | 96 | 65 | 47 | 93 | 31 | 47 | – | 74 | 14 | 18 | 14 |
| BC | 84 | 34 | 21 | 84 | 10 | 13 | 26 | – | 5 | 7 | 10 |
| BCSP | 100 | 92 | 66 | 97 | 62 | 83 | 86 | 95 | – | 63 | 59 |
| BCFP | 98 | 88 | 68 | 98 | 54 | 62 | 82 | 93 | 37 | – | 58 |
| BCDO | 100 | 90 | 79 | 100 | 65 | 67 | 86 | 90 | 41 | 42 | – |
| Algorithm | R | MBP | SH | LLM | SP | FP | DO | BC | BCSP | BCFP | BCDO |
| R | – | 10 | 1 | 41 | 1 | 2 | 5 | 13 | 3 | 1 | 0 |
| MBP | 90 | – | 23 | 93 | 28 | 28 | 32 | 59 | 18 | 11 | 7 |
| SH | 99 | 77 | – | 91 | 48 | 61 | 48 | 73 | 33 | 24 | 25 |
| LLM | 59 | 7 | 9 | – | 1 | 7 | 3 | 14 | 1 | 2 | 2 |
| SP | 99 | 72 | 52 | 99 | – | 48 | 67 | 68 | 35 | 28 | 28 |
| FP | 98 | 72 | 39 | 93 | 52 | – | 64 | 84 | 37 | 29 | 21 |
| DO | 95 | 68 | 52 | 97 | 33 | 36 | – | 63 | 35 | 19 | 20 |
| BC | 87 | 41 | 27 | 86 | 32 | 16 | 37 | – | 10 | 12 | 13 |
| BCSP | 97 | 82 | 67 | 99 | 65 | 63 | 65 | 90 | – | 51 | 40 |
| BCFP | 99 | 89 | 76 | 98 | 72 | 71 | 81 | 88 | 49 | – | 52 |
| BCDO | 100 | 93 | 75 | 98 | 72 | 79 | 80 | 87 | 60 | 48 | – |
| Algorithm | R | MBP | SH | LLM | SP | FP | DO | BC | BCSP | BCFP | BCDO |
| R | – | 11 | 2 | 45 | 5 | 6 | 4 | 11 | 4 | 2 | 2 |
| MBP | 89 | – | 19 | 92 | 36 | 28 | 31 | 60 | 21 | 13 | 22 |
| SH | 98 | 81 | – | 94 | 49 | 57 | 42 | 80 | 33 | 38 | 26 |
| LLM | 55 | 8 | 6 | – | 7 | 1 | 6 | 16 | 5 | 3 | 2 |
| SP | 95 | 64 | 51 | 93 | – | 45 | 54 | 73 | 33 | 31 | 27 |
| FP | 94 | 72 | 43 | 99 | 55 | – | 44 | 67 | 32 | 33 | 39 |
| DO | 96 | 69 | 58 | 94 | 46 | 56 | – | 66 | 28 | 26 | 31 |
| BC | 89 | 40 | 20 | 84 | 27 | 33 | 34 | – | 15 | 10 | 17 |
| BCSP | 96 | 79 | 67 | 95 | 67 | 68 | 72 | 85 | – | 48 | 45 |
| BCFP | 98 | 87 | 62 | 97 | 69 | 67 | 74 | 90 | 52 | – | 49 |
| BCDO | 98 | 78 | 74 | 98 | 73 | 61 | 69 | 83 | 55 | 51 | – |
Appendix D Misc.
| Element Range | Action Description |
| -2 | Default |
| -1 | Forfeit |
| 0 | Pass |
| 1–6 | Switch |
| 7–11 | Move 1 |
| 12–16 | Move 2 |
| 17–21 | Move 3 |
| 22–26 | Move 4 |
| 27–31 | Move 1 + Mega Evolve |
| 32–36 | Move 2 + Mega Evolve |
| 37–41 | Move 3 + Mega Evolve |
| 42–46 | Move 4 + Mega Evolve |
| 47–51 | Move 1 + Z-Move |
| 52–56 | Move 2 + Z-Move |
| 57–61 | Move 3 + Z-Move |
| 62–66 | Move 4 + Z-Move |
| 67–71 | Move 1 + Dynamax |
| 72–76 | Move 2 + Dynamax |
| 77–81 | Move 3 + Dynamax |
| 82–86 | Move 4 + Dynamax |
| 87–91 | Move 1 + Terastallize |
| 92–96 | Move 2 + Terastallize |
| 97–101 | Move 3 + Terastallize |
| 102–106 | Move 4 + Terastallize |
| Component | Specification |
| Learning Rate | 1e-5 |
| Discount Factor () | 1.0 |
| GAE Lambda () | 0.95 |
| Clip Range | 0.2 |
| Entropy Coefficient | 0.001 |
| Value Function Coefficient | 0.5 |
| Max Gradient Norm | 0.5 |
| Number of Steps per Update | 24 * 128 |
| Batch Size | 64 |
| Number of Epochs | 10 |
| Total Timesteps | 5,013,504 |